At what speed v will the length of a 1.00 m stick
Answer #1

L = L0*sqrt[1 - (v/c)^2]

L/L0 = sqrt[1 - (v/c)^2]

(L/L0)^2 = 1 - (v/c)^2

(v/c)^2 = 1 - (L/L0)^2

v/c = sqrt[1 - (L/L0)^2]

In this case L/L0 = 0.4

v/c = sqrt[1 - (0.4)^2]

v/c = sqrt(0.84)

v = 0.9165c